# DuoCast: Duo-Probabilistic Diffusion for Precipitation Nowcasting

### v3

@@ Introduction
To this end, we next we investigate two architectural designs that (approximately) satisfy the conditions of Corollary 3, corresponding to the low-frequency and high-frequency diffusion models.
Low-Frequency Diffusion Model
Weather fronts, which evolve gradually and span broad spatial regions, are key to shaping precipitation systems (Catto et al. 2012). For instance, warm fronts bring steady light rain, cold fronts induce abrupt heavy rainfall, and occluded fronts result in widespread precipitation. Inadequate modeling of these fronts can lead to significant forecasting errors. As shown in Fig. 1(a), the context indicates a weakening warm front, yet DiffCast erroneously predicts intensifying rainfall. Therefore, the low-frequency diffusion captures these trends with convolutions to guide prediction, conditioned on front dynamics derived from underlying air mass interactions.
Air Mass Spatial Modeling.
To model the spatial organization of weather fronts, we introduce a convolutional module that learns the distribution and interactions of underlying air masses. As air mass boundaries often coincide with sharp gradients in precipitation (Lagerquist, Mcgovern, and Gagne II 2019), this module extracts spatial cues relevant to front structure. Specifically, we apply three tailored convolutional layers to the preliminary forecasts $\bar{\mathbf{Y}}$ and $\mathbf{Y}^{\dagger}$ to capture local context, multi-scale structure, and temporal anchoring, forming a spatial representation $\mathbf{A}$.
$\mathbf{A}=\text{Concat}(\bar{\mathbf{A}},\mathbf{A}^{\dagger}).$
Weather Front Temporal Modeling.
To capture the temporal evolution of weather fronts, we devise a lightweight convolution-based cross-attention mechanism that relates the current spatial features to both historical structure and recent trends. Specifically, for each frame $\mathbf{A}_{n}$, we compute attention over two reference frames: (1) the leading frame $\mathbf{A}_{1}$, which serves as a static structural prior, and (2) the previous frame $\mathbf{A}_{n-1}$, which reflects recent changes in air mass interactions. The attention query is derived from the current frame, while the key and value come from the concatenation of the leading and previous frames. Formally: Cross-Attention($\mathbf{Q}=\mathbf{W}_{\mathbf{Q}}\circledast\mathbf{A}_{n}$$\mathbf{K}=\mathbf{W}_{\mathbf{K}}\circledast\left[\mathbf{A}_{1}\oplus\mathbf{A}_{n-1}\right]$$\mathbf{V}=\mathbf{W}_{\mathbf{V}}\circledast\left[\mathbf{A}_{1}\oplus\mathbf{A}_{n-1}\right]$),
[…]
High-Frequency Diffusion Model
Optimization
$\mathcal{L}_{2}=\lambda_{1}\mathcal{L}_{P}+\lambda_{2}\mathcal{L}_{\text{low}}+\lambda_{3}\mathcal{L}_{\text{high}},$ (10)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are hyperparameters.
Experiments & Discussions
Experimental Settings
Datasets
To evaluate DuoCast’s effectiveness in generating precise precipitation maps, we performed experiments using four radar echo datasets: SEVIR (Veillette, Samsi, and Mattioli 2020), MeteoNet (Larvor and Berthomier 2021), Shanghai_Radar (Chen et al. 2020), and CIKM [^1] https://tianchi.aliyun.com/dataset/1085. More details regarding datasets are provided in Supplementary.
Evaluation Metrics
Following (Yu et al. 2024; Lin et al. 2025; Gong et al. 2024), we evaluate the nowcasting with the average Critical Success Index (CSI), Heidke Skill Score (HSS), Structural Similarity Index Measure (SSIM), and MSE. The CSI, similar to the Intersection over Union (IoU), measures the degree of pixel-wise alignment between predictions and ground truth after thresholding them into binary (0/1) matrices, while HSS reflects improvement over random guessing. Following (Lin et al. 2025), we measure and average CSI (CSI-M) and CSI with two highest thresholds to measure the performance of high-intensity precipitation forecasting. SSIM and MSE assess prediction quality.
Implementation Details
and performance regarding hyper-parameters are provided in supplementary materials.
@@
DuoCast 0.338 0.182 0.107 0.432 0.683 463.1
High-Freq 0.310 0.158 0.050 0.401 0.611 470.4
Low-Freq 0.327 0.164 0.066 0.410 0.642 462.0
Low-Frequency Backbone Exploration
Attention 0.310 0.143 0.043 0.393 0.608 562.5
Conv 0.327 0.164 0.066 0.410 0.642 462.0
- $\text{Conv}_{\text{I}}$ 0.316 0.155 0.051 0.405 0.631 470.1
- $\text{Conv}_{\text{II}}$ 0.317 0.159 0.048 0.400 0.632 472.0
- $\text{Conv}_{\text{III}}$ 0.320 0.156 0.050 0.404 0.622 480.6
- $\mathbf{Y^{\dagger}}$ 0.316 0.146 0.048 0.400 0.619 490.7
- Airmass 0.309 0.143 0.051 0.388 0.625 492.2
- Front 0.311 0.138 0.050 0.383 0.628 471.3
Compared with the State-of-the-Art
To assess the performance of our DuoCast framework in precipitation nowcasting, we compare it against a range of state-of-the-art methods, including probabilistic, deterministic, and hybrid approaches: MAU (Chang et al. 2021), SimVP (Gao et al. 2022b), FourCastNet (Pathak et al. 2022), Earthformer (Gao et al. 2022a), PhyDnet (Guen and Thome 2020), EarthFarseer (Wu et al. 2024), NowcastNet (Zhang et al. 2023), PreDiff (Gao et al. 2024), DiffCast (Yu et al. 2024), CasCast (Gong et al. 2024), FACL (Yan et al. 2024), and AlphaPre (Palmer et al. 2005).
Method SEVIR MeteoNet
↑CSI-M ↑CSI-181 ↑CSI-219 ↑HSS ↑SSIM ↓MSE ↑CSI-M ↑CSI-24 ↑CSI-32 ↑HSS ↑SSIM ↓MSE
MAU (2021) 0.3076 0.1071 0.0516 0.3863 0.6505 355.48 0.3233 0.2839 0.0997 0.4452 0.7897 12.92
SimVP (2022) 0.3108 0.1106 0.0517 0.3924 0.6508 383.56 0.3351 0.3002 0.1130 0.4573 0.7804 13.45
FourCastNet (2022) 0.2686 0.0717 0.0339 0.3355 0.5976 410.27 0.3027 0.2533 0.1085 0.4216 0.6450 15.05
Earthformer (2022) 0.2892 0.0844 0.0245 0.3665 0.6633 360.11 0.3205 0.2884 0.1237 0.4491 0.7772 14.43
PhyDNet (2020) 0.3017 0.1040 0.0278 0.3812 0.6532 357.63 0.3384 0.3194 0.1366 0.4673 0.7823 14.48
EarthFarseer (2024) 0.3004 0.0992 0.0413 0.3829 0.6327 388.91 0.3404 0.3170 0.1372 0.4726 0.7542 14.10
NowcastNet (2023) 0.2791 0.0770 0.0351 0.3512 0.6839 412.94 0.3427 0.3206 0.1598 0.4751 0.7879 15.64
PreDiff (2024) 0.2744 0.0627 0.0235 0.3592 0.5881 409.52 0.2944 0.2641 0.1202 0.4462 0.7085 17.95
DiffCast (2024) 0.3050 0.1300 0.0582 0.3996 0.6482 559.59 0.3512 0.3340 0.1808 0.4846 0.7887 17.93
CasCast (2024) 0.2878 0.0954 0.0312 0.3563 0.5680 471.14 0.3317 0.3021 0.1351 0.4682 0.7622 15.13
FACL (2024) 0.3161 0.1349 0.0655 0.4033 0.5686 417.23 0.3612 0.3341 0.1801 0.5006 0.7881 13.42
AlphaPre (2025) 0.3259 0.1332 0.0545 0.4110 0.6884 345.18 0.3824 0.3633 0.2002 0.5164 0.7968 12.74
Ours 0.3375 0.1818 0.1074 0.4318 0.6827 463.07 0.3892 0.3841 0.2381 0.5297 0.7981 15.82
Method Shanghai Radar CIKM
↑CSI-M ↑CSI-35 ↑CSI-40 ↑HSS ↑SSIM ↓MSE ↑CSI-M ↑CSI-35 ↑CSI-40 ↑HSS ↑SSIM ↓MSE
MAU (2021) 0.3983 0.3621 0.2417 0.5346 0.7195 30.40 0.3039 0.2054 0.1241 0.3928 0.6325 40.74
SimVP (2022) 0.3850 0.3549 0.2382 0.5194 0.7795 34.40 0.3052 0.2044 0.1321 0.3955 0.6538 38.06
FourCastNet (2022) 0.3571 0.3108 0.2073 0.4868 0.5598 32.10 0.2980 0.1849 0.1015 0.3801 0.4359 36.14
Earthformer (2022) 0.3503 0.3178 0.1872 0.4844 0.7298 35.57 0.3077 0.2039 0.1369 0.4001 0.6267 36.49
PhyDNet (2020) 0.3654 0.3236 0.2176 0.4957 0.7751 36.41 0.3038 0.2052 0.1287 0.3931 0.6541 39.56
EarthFarseer (2024) 0.3926 0.3608 0.2343 0.5330 0.5405 32.68 0.3000 0.2046 0.1259 0.3911 0.6373 39.87
NowcastNet (2023) 0.3953 0.3608 0.2450 0.5334 0.7902 33.56 0.2991 0.1940 0.1188 0.3865 0.6713 44.46
PreDiff (2024) 0.3504 0.3026 0.1855 0.4991 0.6981 40.56 0.2842 0.1848 0.1002 0.3481 0.5582 17.95
DiffCast (2024) 0.4089 0.3740 0.2606 0.5476 0.7879 36.35 0.3159 0.2009 0.1457 0.4085 0.6499 42.78
CasCast (2024) 0.3651 0.3291 0.2192 0.4971 0.7250 34.51 0.3021 0.1974 0.1142 0.3806 0.5241 46.14
FACL (2024) 0.3940 0.3511 0.2355 0.5308 0.7601 34.62 0.3105 0.1993 0.1354 0.3996 0.6442 40.72
AlphaPre (2025) 0.4178 0.3854 0.2615 0.5534 0.7951 28.02 0.3194 0.2068 0.1416 0.4137 0.6568 35.18
Ours 0.4252 0.3933 0.2821 0.5643 0.7788 30.20 0.3179 0.2275 0.1562 0.4217 0.6659 36.10
Quantitative Analysis
Qualitative Analysis
Figure 3: Qualitative comparison with SoTA on a SEVIR event. DuoCast captures finer micro-scale details (pink box) and preserves more consistent evolution trends (red line).
Ablation Study
Effectiveness of Low-Frequency Diffusion Model.
Effectiveness of High-Frequency Diffusion Model.
Effectiveness of Air Mass Modeling.
As shown in Table 1, removing air mass spatial modeling leads to a notable drop in CSI, particularly CSI-M, indicating its importance in capturing the spatial evolution of weather fronts. Additionally, we analyze the effectness of convolution in air mass spatial modeling. As shown in the Fig. 5, $\text{Conv}_{\text{I}}$, captures the spatial dynamics of surrounding air mass effects, highlighted by the heatmap. The frontal convolution, $\text{Conv}_{\text{II}}$, focuses on occluded fronts, which occur in critical transition areas where cold and warm air masses interact, as depicted by the figure’s red band. The temporal convection convolution, $\text{Conv}_{\text{III}}$, captures pixel-level temporal convection, with the red areas indicating potential temporal motion. Quantitatively, as shown in Table 1, every component in air mass spatial modeling contributes to superior performance.
Effectiveness of Weather Front Modeling.
Effectiveness of $\mathbf{Y^{\dagger}}$ for High-Intensity Precipitation
To assess the role of $\mathbf{Y^{\dagger}}$ in modeling high-intensity precipitation, we conduct targeted experiments on the SEVIR dataset. As shown in Table 1, removing $\mathbf{Y^{\dagger}}$ results in substantial performance drops for severe rainfall events, with CSI-181 and CSI-219 decreasing by 10.8% and 27.0%, respectively. Figure 4(d) visualizes high-intensity events (values $\geq$ 181), where predictions with $\mathbf{Y^{\dagger}}$ align closely with the ground truth, consistently maintaining the precipitation band over time. In contrast, models without $\mathbf{Y^{\dagger}}$ increasingly miss key regions, and the predicted band degrades notably after 30 minutes. These results highlight the importance of $\mathbf{Y^{\dagger}}$ in learning temporal consistency and structure in severe precipitation. Furthermore, Figure 6 (b) shows that our model achieves consistently higher CSI across thresholds, confirming its advantage in capturing high-intensity rainfall.
Figure 6: CSI across different prediction time steps and intensity thresholds for various methods on the SEVIR dataset.
Conclusion
In this paper, we introduce DuoCast - a dual-probabilistic meteorology-aware model designed for preservating different frequency information. Specially, low-frequency aims to capture the low frequency broader trend of weather patterns and high-frequency is designed to capture high frequency micro-scale variability relevant to precipitation. Comprehensive experiments on four real-world datasets validate the effectiveness of the proposed framework.

### v4

@@ Appendix B Formal Proofs
Proof.
Apply integration by parts along each coordinate axis. ∎
@@
Lemma 3 (Tail energy under BV envelope).
Let $f=\mathbf{Y^{0}_{\text{low}}}\in L^{2}(\mathbb{R}^{d})$ and denote $g:=H_{L}*f$. For any radius $\Omega>0$,
$\int_{|\xi|>\Omega}|\widehat{g}(\xi)|^{2}\,d\xi\leq C^{2L}\!\int_{|\xi|>\Omega}(1+|\xi|)^{-2L}\,|\widehat{f}(\xi)|^{2}\,d\xi.$
If $L>d/2$, the integral on the right is finite.
Proof.
Since $\widehat{g}=H_{L}\,\widehat{f}$ and $|H_{L}(\xi)|\leq C^{L}(1+|\xi|)^{-L}$, the claim follows by direct substitution. ∎
Corollary 3 ($\varepsilon$-tail radius).
Fix $\varepsilon\in(0,1)$. Any radius $\Omega_{\varepsilon}^{\sharp}(L)$ satisfying
$C^{2L}\!\int_{|\xi|>\Omega_{\varepsilon}^{\sharp}(L)}(1+|\xi|)^{-2L}\,|\widehat{f}(\xi)|^{2}\,d\xi\leq\varepsilon\,\|f\|_{2}^{2}$
ensures $\int_{|\xi|>\Omega_{\varepsilon}^{\sharp}(L)}|\widehat{g}(\xi)|^{2}\leq\varepsilon\,\|f\|_{2}^{2}.$
Definition 1 (Input–aware $\varepsilon$ cut‑off).
For $f\in L^{2}(\mathbb{R}^{d})$ and $\varepsilon\in(0,1)$ define $\Omega_{\varepsilon}^{\sharp}(L)$ as:
$\inf\Bigl\{\Omega>0:C^{2L}\!\!\int_{|\xi|>\Omega}(1+|\xi|)^{-2L}\,|\widehat{f}(\xi)|^{2}\,d\xi\leq\varepsilon\,\|f\|_{2}^{2}\Bigr\}.$
By Lemma 3, any feature $g=H_{L}*f$ then satisfies $\int_{|\xi|>\Omega_{\varepsilon}^{\sharp}(L)}|\widehat{g}(\xi)|^{2}\leq\varepsilon\,\|f\|_{2}^{2}.$
@@
Let $\mathbf{Y}\in L^{2}(\mathbb{R}^{d})$ be any target, $\omega_{c}=\Omega_{\varepsilon}^{\sharp}(L)$, $\mathbf{Y}_{\text{low}}=P_{\mathcal{B}_{\omega_{c}}}(\mathbf{Y})$ and $\mathbf{Y}_{\text{high}}=P_{\mathcal{B}_{\omega_{c}}^{\perp}}(\mathbf{Y})=\mathbf{Y}-\mathbf{Y}_{\text{low}}$. For every generation $g=\mathbf{Y}^{t}$ based on a CNN with input $f=\mathbf{Y}^{0}$,
@@
Proof.
Write the orthogonal decompositions $\mathbf{Y}^{t}=\mathbf{Y}^{t}_{\text{Low}}+\mathbf{Y}^{t}_{\text{High}}$ and $\mathbf{Y}=\mathbf{Y}_{\text{Low}}+\mathbf{Y}_{\text{High}}$ with $\mathbf{Y}^{t}_{\text{low}}=P_{\mathcal{B}_{\omega_{c}}}(\mathbf{Y}^{t})$ and $\mathbf{Y}^{t}_{\text{high}}=P_{\mathcal{B}_{\omega_{c}}^{\perp}}(\mathbf{Y}^{t})$. By Definition 1,
$\|\mathbf{Y}^{t}_{\text{High}}\|_{2}^{2}\;=\;\!\int_{|\xi|>\omega_{c}}|\widehat{g}(\xi)|^{2}\,d\xi\;\leq\;\varepsilon\,\|f\|_{2}^{2}.$
Parseval and projection orthogonality give
$\|\mathbf{Y}-\mathbf{Y}^{t}\|_{2}^{2}\;=\;\|\mathbf{Y}_{\text{High}}-\mathbf{Y}^{t}_{\text{High}}\|_{2}^{2}+\|\mathbf{Y}_{\text{Low}}-\mathbf{Y}^{t}_{\text{Low}}\|_{2}^{2}.$
Drop the first (non–negative) term:
$\|\mathbf{Y}-\mathbf{Y}^{t}\|_{2}^{2}\;\geq\;\|\mathbf{Y}_{\text{High}}-\mathbf{Y}^{t}_{\text{High}}\|_{2}^{2}.$
Reverse triangle inequality in $L^{2}$:
$\|\mathbf{Y}_{\text{High}}-\mathbf{Y}^{t}_{\text{High}}\|_{2}\;\geq\;\bigl|\|\mathbf{Y}_{\text{High}}\|_{2}-\|\mathbf{Y}_{\text{High}}^{t}\|_{2}\bigr|.$
Combine with $\|g_{\mathrm{H}}\|_{2}\leq\sqrt{\varepsilon}\|f\|_{2}$, square both sides and obtain
$\|\mathbf{Y}_{\text{High}}-\mathbf{Y}^{t}_{\text{High}}\|_{2}\;\geq\;\Bigl[\|\mathbf{Y}_{\text{High}}\|_{2}-\sqrt{\varepsilon}\,\|\mathbf{Y}^{0}\|_{2}\Bigr]_{+}.$
@@
Proof.
Fix an arbitrary $\delta>0$. By density of $\mathcal{G}_{\theta}$, choose $\theta$ such that $\lVert\mathbf{Y}_{\text{Low}}-g^{(1)}_{\theta}\rVert_{2}<\delta$. Set $e_{\text{Low}}:=\mathbf{Y}_{\text{Low}}-g^{(1)}_{\theta}$. Similarly, density of $\mathcal{G}_{\phi}$ provides $\phi$ with $\lVert\mathbf{Y}_{\text{High}}-g^{(2)}_{\phi}\rVert_{2}<\delta$; define $e_{\text{High}}:=\mathbf{Y}_{\text{High}}-g^{(2)}_{\phi}$. Because $e_{\mathrm{L}}\in\mathcal{B}_{\omega_{c}}$ and $e_{\mathrm{H}}\in\mathcal{B}_{\omega_{c}}^{\perp}$ are orthogonal,
$\lVert\mathbf{Y}-g^{(1)}_{\theta}-g^{(2)}_{\phi}\rVert_{2}^{2}=\lVert e_{\mathrm{L}}\rVert_{2}^{2}+\lVert e_{\mathrm{H}}\rVert_{2}^{2}<\delta^{2}+\delta^{2}=2\delta^{2}.$
Letting $\delta\to 0$ proves the claim. ∎
Method SEVIR
↑CSI ↑CSI-181 ↑CSI-219 ↑HSS ↑SSIM ↓MSE
$\lambda_{1}$ = 0.5 , $\lambda_{2}$ = 0.1 , $\lambda_{3}$ = 0.1 0.3255 0.1712 0.0963 0.4263 0.6753 480.83
$\lambda_{1}$ = 0.1 , $\lambda_{2}$ = 0.5 , $\lambda_{3}$ = 0.1 0.3247 0.1754 0.0925 0.4048 0.6682 490.13
$\lambda_{1}$ = 0.1 , $\lambda_{2}$ = 0.1 , $\lambda_{3}$ = 0.5 0.3375 0.1818 0.1074 0.4318 0.6827 463.07
Table 4: Effect of loss weight hyper-parameters in Eq. (10) on SEVIR for the DuoCast model.
@@
w/o $\mathbf{A}_{1}$ 0.320 0.157 0.060 0.398 0.633 465.4
w/ $\mathbf{A}_{1}$ 0.327 0.164 0.066 0.410 0.642 462.0
Table 5: Ablation study showing the effectiveness of $\mathbf{A}_{1}$.
Appendix C Datasets Details
The effectiveness of DuoCast was evaluated on four widely used radar echo datasets: SEVIR, MeteoNet, Shanghai_Radar and CIKM [^2] https://tianchi.aliyun.com/dataset/1085.
SEVIR captures both storm and random events in the United States from 2017 to 2019. It contains 20,393 sequences of radar frames representing weather events, each covering a 4-hour period with a spatial coverage of $\times384 km$. Every pixel corresponds to an area of $\times1 km$, and the data has a 5-minute temporal resolution.
MeteoNet contains rain radar data from the northwest and southeast regions of France, covering the period from 2016 to 2018. The radar data in MeteoNet has a spatial resolution of $0.01^{\circ}$, with observations recorded at a 6-minute temporal resolution. Following SoTA methods, we used radar observations specifically from the northwest region of France.
Shanghai_Radar consists of continuous radar echo frames generated by volume scans at an approximately 6-minute temporal resolution, collected from October 2015 to July 2018 in Pudong, Shanghai. Each radar echo map covers an area of $\times501 km$.
CIKM is from the CIKM AnalytiCup 2017 Competition, as a radar dataset that records precipitation samples over an $\times101 km$ area in Guangdong, China. Each sequence includes 15 radar echo maps as a sample, with a 6-minute temporal resolution.
We follow the experimental protocols of state-of-the-art methods, such as DiffCast and AlphaPre, across all datasets, maintaining the original temporal resolution while reducing the spatial dimensions to $128\times 128$. Additionally, since our primary focus is on modeling precipitation events, we segment continuous sequences into multiple events for the SEVIR, MeteoNet, Shanghai_Radar, and CIKM datasets, exactly based on the procedure used in prior SOTA method. For all datasets except CIKM, we predict 20 future frames given 5 past frames ($5\rightarrow 20$, i.e., $L_{\text{in}}=5$, $L_{\text{out}}=20$). For CIKM, due to its shorter sequence length, we predict only 10 future frames given 5 past frames ($5\rightarrow 10$, i.e., $L_{\text{in}}=5$, $L_{\text{out}}=10$).
Dataset $N_{tr}$ $N_{va}$ $N_{te}$ $(C,H,W)$ $T_{i}$ $T_{o}$
SEVIR 23808 6016 8100 $(1,128,128)$ 5 20
MeteoNet 6308 1310 1310 $(1,128,128)$ 5 20
Shanghai 1534 526 526 $(1,128,128)$ 5 20
CIKM 8000 2000 4000 $(1,128,128)$ 5 10
Table 6: Dataset statistics. $N_{tr}$, $N_{va}$ and $N_{te}$ represent the number of samples in the training, validation, and test sets, respectively.
Appendix D Implementation Details
We trained the DuoCast framework using the Adam optimizer with a learning rate of 0.0001. In line with standard practices for diffusion models, we used 1,000 diffusion steps for both the low- and high-frequency branches. The batch sizes were set to 4 for the low-frequency model and 3 for the high-frequency model. We adopt a two-stage training strategy. In the first stage, only the low-frequency model is trained. In the second stage, both the low- and high-frequency models are trained jointly. The learning rate is kept the same in both stages. However, we apply loss weighting via $\lambda$ to stabilize training. In the first stage, the loss weight coefficients were empirically set to $\lambda_{1}=\lambda_{2}=0.5$. In the second stage, the loss weight coefficients were empirically set to $\lambda_{1}=\lambda_{2}=0.1$ and $\lambda_{3}=0.5$ without further tuning. The SimVP model is used as initial predictor $P$. The cutoff $\theta_{\text{init}}$ is set as the high-intensity precipitation threshold defined in the datasets SEVIR dataset $\theta_{\text{init}}$=181, MeteoNet dataset $\theta_{\text{init}}$=24, Shanghai and CIKM $\theta_{\text{init}}$=35.
Appendix E Additional Analysis
In this section, we provide an extra analysis on the design of framework loss, the computational complexity and the effectiveness of our model design.
Hyper-Parameters for Loss Weights. Starting from an empirical setting ($\lambda_{1}=\lambda_{2}=0.1$ and $\lambda_{3}=0.5$), We conducted sensitivity experiments (Table 4) to assess the impact of different weight configurations. The results indicate that the model performs robustly across a range of reasonable settings, with the combination $(0.1,0.1,0.5)$ achieving the best overall performance across all metrics. In contrast, the $(0.5,0.1,0.1)$ setting assigns equal emphasis to the low- and high-frequency models, causing the initially well-trained low-frequency model to become significantly biased as it adjusts to fit the high-frequency component during subsequent training. Similarly, the $(0.1,0.5,0.1)$ setting gives excessive weight to the low-frequency model, leading to over-optimization of the low-frequency component while leaving the high-frequency model under-trained.
Figure 7: Qualitative comparison between w/o and w/ $\mathbf{A}_{1}$ on SEVIR.
Model Complexity. In Table 7, we compare the number of parameters and TFLOPs of our method and existing baselines, evaluated on a single NVIDIA RTX 4090 GPU using the SEVIR dataset. DuoCast has more parameters than DiffCast but requires fewer TFLOPs, indicating improved computational efficiency. Compared to AlphaPre, DuoCast has a similar parameter count but incurs higher computational cost in terms of TFLOPs. However, given the significant advances in GPU hardware—such as the NVIDIA RTX 4090, which supports up to 82.6 TFLOPs (FP32) and 330.4 TFLOPs (FP16), DuoCast’s demand of 42.16 TFLOPs remains well within the capabilities of modern hardware. Considering the importance of inference speed in precipitation nowcasting, DuoCast achieves a practical balance between performance and efficiency.
Additionally, DuoCast is efficient, with inference taking 0.56 seconds per sample (1 sample = 20 frames prediction; 100-minute horizon for SEVIR) on an NVIDIA RTX 4090, while the time of DiffCast is 3.67 seconds. This latency is suitable for running precipitation nowcasting smoothly on consumer GPUs.
Method Complexity
# Param TFLOPs Year
MCVD 105.91M 4.69 2022
Prediff 135.24M 2.80 2023
STRPM 439.63 M 0.28 2022
SimVP 44.25M 0.05 2022
Earthformer 34.61M 0.04 2022
MAU 20.13M 0.09 2021
ConvGRU 18.21M 0.03 2017
PhyDnet 11.80M 0.08 2020
DiffCast 58.33M 72.49 2024
FACL 14.38M 0.02 2024
AlphaPre 89.03M 1.56 2025
DuoCast 94.18M 0.30 2025
Table 7: Analysis of model complexity with SOTA methods.
Effectiveness of $\mathbf{A}_{1}$ for Temporal Modeling. The gap between $\mathbf{A}_{1}$ and $\mathbf{A}_{n-1}$ captures essential temporal dynamics. Excluding $\mathbf{A}_{1}$ removes the model’s global temporal anchor, making it difficult to maintain coherent long-range structure. As a result, low-frequency components begin to drift over time, leading to distorted large-scale patterns and reduced forecast stability. In Fig. 7, without $\mathbf{A}_{1}$, the prediction aligns closely with the previous frame but fails to identify the change trend. For instance, in the 100-minute frame, though the high-intensity precipitation band aligns well with the preceding frame, its shape becomes distorted and diverges significantly from that in the first frame. The quantitative results in Table 6 further support this observation.
The Statistical Significance of Improvement. To assess statistical significance, we conducted paired sample t-tests between DuoCast and AlphaPre on the SEVIR dataset (CSI) across 8100 samples. Table 8 shows significant improvements. This confirms that DuoCast’s improvements are statistically significant.
CSI-Threshold p-value DuoCast >> AlphaPre
M 0.035 Significant difference
181 <<0.001 Significant difference
219 0.001 Significant difference
Table 8: The paired sample t-tests between DuoCast and AlphaPre on the SEVIR dataset (CSI).
Appendix F More Qualitative Results
Additional qualitative results are shown in Figures 811 for the four datasets. Compared to the most recent method AlphaPre, DuoCast generate sharper predictions and more accurately captures high-intensity precipitation regions, particularly in long lead-time forecasts. When compared to the baseline model DiffCast, both models perform similarly well in early frames, effectively capturing overall trends and fine-grained details. However, DiffCast’s predictions degrade in later frames, exhibiting spatial distortions and loss of structural consistency, whereas DuoCast preserves both detail and stability over extended time horizons.
Figure 8: Prediction examples on the SEVIR dataset.
Figure 9: Prediction examples on the MeteoNet dataset.
Figure 10: Prediction examples on the Shanghai dataset.
Figure 11: Prediction examples on the CIKM dataset.